OpenStudy (josee):

lim as x goes to 2 of (f(x))-f(2))/(x-2) when f(x)= 5x^2-x+1

OpenStudy (anonymous):

19

OpenStudy (josee):

what did you do? Did you plug in the equation for the top part and pug in 2 for f(2)?

OpenStudy (anonymous):

f(x)-f(2)=5x^2-x+1-(5(2^2)-2+1)=5x^2-x+1-19 =5x^2-x-18=(x-2)(5x+9) so thats whats on top so u have: \[\frac{(x-2)(5x+9)}{x-2}\]

OpenStudy (anonymous):

cancel then solve for x=2

OpenStudy (anonymous):

numerator simplifies into (5x²-x-18), that can be factorized into (x-2)(5x+9),, simplify

OpenStudy (anonymous):

the limit=18 by l'hopitals rule

OpenStudy (josee):

it's 19

OpenStudy (anonymous):

no, take the derivative of the numerator and the derivative of the denominator and then plug in 2. you get 18

OpenStudy (anonymous):

by l'hopital's rule:\[\lim_{x \rightarrow 2} \frac{\frac{d}{dx}(f(x)-f(2)}{\frac{d}{dx}(x-2)}=\frac{f \prime(x)-0}{1-0}=f \prime (x)= f \prime (2)\]

OpenStudy (anonymous):

the problem is basically asking for the slope at f(2). So to determine slopes, you calculate the first derivative at that point. Thus, the derivative, f'(x)=10x-2. So f'(2)=18

OpenStudy (anonymous):

derivative= 10x-1 10*2-1=20-1=19

OpenStudy (anonymous):

you are right, haha good call